# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ III-B Communication Modes
For convenience, we introduce the inner product between vector functions $\mathsf{A}({\bf{r}})$ and $\mathsf{B}({\bf{r}})$, defined on the generic space ${\cal{S}}$, as
${\mathsf{A}({\bf{r}})}\,,{\mathsf{B}({\bf{r}})}\right>=\int_{{\cal{S}}}\mathsf{A}({\bf{r}})\bm{\cdot}\mathsf{B}^{*}({\bf{r}})\,\text{d}{\bf{r}}\,.$ (10)
Two vector functions are said to be orthogonal if ${\mathsf{A}({\bf{r}})}\,,{\mathsf{B}({\bf{r}})}\right>=0$.
@@
A fundamental question is determining how many orthogonal communication channels—referred to as communication modes—can be exploited between a source space $\mathcal{S}_{\text{T}}$ and a receiver space $\mathcal{S}_{\text{R}}$ while maximizing the coupling intensity (see Fig. 2). This is associated with the optimal approximation of every element in the image space of a Hilbert-Schmidt operator in terms of singular functions. Specifically, define $\mathcal{X}={\mathbb{L}}^{2}(\mathcal{S}_{\text{T}})$ and $\mathcal{Y}={\mathbb{L}}^{2}(\mathcal{S}_{\text{R}})$ the Hilbert spaces corresponding to the square-integrable functions defined in $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$, respectively. The function ${\mathsf{E}}({\bf{r}})\in\mathcal{Y}$ can be seen as the image of ${\mathsf{J}}({\bf{r}})\in\mathcal{X}$ through the Hilbert-Schmidt kernel $\mathsf{G}\left({\bf{r}},{\bf{s}}\right)=\mathsf{G}\left({\bf{r}}-{\bf{s}}\right)$ on $\mathcal{S}_{\text{T}}\times\mathcal{S}_{\text{R}}$, which induces the operator $\mathbb{G}:\mathcal{X}\rightarrow\mathcal{Y}$ such that, for any ${\mathsf{J}}({\bf{r}})\in\mathcal{X}$, we can write (1). Define the following self-adjoint Hilbert-Schmidt operators $\mathbb{G}\,\mathbb{G}^{*}$ and $\mathbb{G}^{*}\,\mathbb{G}$ and let $\{{\bm{\varphi}}_{n}({\bf{r}})\}$ and $\{{\mathbf{\Psi}}_{n}({\bf{r}})\}$ be the eigenfunctions, associated with eigenvalues $\sigma_{n}^{2}$, of the following coupled eigenfunction problems [38]
$\left(\mathbb{G}\,\mathbb{G}^{*}\,{\bm{\varphi}}_{n}\right)({\bf{r}})=\sigma_{n}^{2}\,{\bm{\varphi}}_{n}({\bf{r}})\quad\quad\quad\quad\quad\quad\left(\mathbb{G}^{*}\,\mathbb{G}\,{\mathbf{\Psi}}_{n}\right)({\bf{r}})=\sigma_{n}^{2}\,{\mathbf{\Psi}}_{n}({\bf{r}})\,.$ (11)
@@
An important property of Hilbert-Schmidt operators is that they are compact and admit either a finite or countably infinite orthonormal basis. In particular, two sets of orthonormal eigenfunctions $\{{\bm{\varphi}}_{n}({\bf{r}})\}$, $\{{\mathbf{\Psi}}_{n}({\bf{r}})\}$ exist, which are solutions, respectively, of the coupled eigenfunction problems in (11) with the same real eigenvalues $\sigma_{1}^{2}\geq\sigma_{2}^{2}\geq\sigma_{3}^{2}\ldots$ [38, 39]. Note that $\left\{{\bm{\varphi}}_{n}({\bf{r}})\right\}$ and $\left\{{\mathbf{\Psi}}_{n}({\bf{r}})\right\}$ are two sets of orthonormal (vector) functions that are complete, respectively, in $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$. The orthonormal condition implies that ${{\bm{\varphi}}_{k}({\bf{r}})}\,,{{\bm{\varphi}}_{n}({\bf{r}})}\right>=\delta_{k-n}$, and similarly, the same condition holds for the set $\{{\mathbf{\Psi}}_{n}({\bf{r}})\}$. In general, the eigenvalues have significant magnitudes up to a certain value $\sigma_{N}^{2}$, i.e., the sum of the $\sigma_{n}^{2}$’s for $n>N$ is negligible with respect to the sum of the first $N$ eigenvalues, so that we can consider up to $N$ well-coupled communication modes between the two spaces. The number $N$ is often called the number of DoF available in the link. Detailed discussion on optimal approximation and the definition of DoF can be found in [38].
@@
In summary, the finite coupling operator $\mathbb{G}$ between finite spaces $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$ is a Hilbert-Schmidt operator, and up to $N$ communication modes (i.e., orthogonal channels) can be identified for a given level of approximation accuracy, each of them represented by the triad $\left({\bm{\varphi}}_{n}({\bf{r}}),{\mathbf{\Psi}}_{n}({\bf{r}}),\sigma_{n}\right)$. It follows that the current density ${\mathsf{J}}({\bf{r}})$ on the space $\mathcal{S}_{\text{T}}$, i.e., for ${\bf{r}}\in\mathcal{S}_{\text{T}}$, can be well approximated as a linear combination of the basis functions
${\mathsf{J}}({\bf{r}})$ $\simeq\sum_{n=1}^{N}j_{n}\,{\bm{\varphi}}_{n}({\bf{r}})$ (12)
@@
where the complex coefficients $j_{n}$ are given by $j_{n}=\left<{\mathsf{J}}({\bf{r}}),{\bm{\varphi}}_{n}({\bf{r}})\right>$, for $n=1,2,\ldots,N$ Similarly, it holds for ${\mathsf{E}}({\bf{r}})$ in $\mathcal{S}_{\text{R}}$ through the basis set $\left\{{\mathbf{\Psi}}_{n}({\bf{r}})\right\}$
${\mathsf{E}}({\bf{r}})$ $\simeq\sum_{n=1}^{N}e_{n}\,{\mathbf{\Psi}}_{n}({\bf{r}})=\sum_{n=1}^{N}\sigma_{n}\,j_{n}\,{\mathbf{\Psi}}_{n}({\bf{r}})\,.$ (13)
@@
Figure 3: A source and receiver space interacting through a reconfigurable scattering device.
@@ III-C Reconfigurable Scattering Device: Mode Transfer Matrix
Consider now the scenario illustrated in Fig. 3, where a source space $\mathcal{S}_{\text{T}}$ (e.g., a transmitter) and a receiver space $\mathcal{S}_{\text{R}}$ (the receiver) interact through a reconfigurable scattering device. We consider the reconfigurable scattering device to be enclosed within the finite space ${\cal{S}}$. The impinging EM field generated by currents located in the source space $\mathcal{S}_{\text{T}}$ induces a current distribution ${\mathsf{J}}_{\mathrm{s}}({\bf{r}})$ on the scattering device, with ${\bf{r}}\in{\cal{S}}$, that radiates the scattered field ${\mathsf{E}}_{\mathrm{s}}({\bf{r}})$ [35]. [^3] A more general treatment is based on the equivalent principle where fictitious electric and magnetic currents are introduced as a function of the tangent electric and magnetic fields [15]. For clarity, hereby we focus on a more limited case valid for dielectric and non-perfect conductors.
For any linear time-invariant scattering object, the induced current density ${\mathsf{J}}_{\mathrm{s}}({\bf{r}})$, ${\bf{r}}\in{\cal{S}}$, is a linear functional of the EM field ${\mathsf{E}}({\bf{r}})$ that can be described as
@@
where $\mathsf{D}({\bf{r}},{\bf{s}})$ is the impulse response dyadic that fully describes the relationship between the electric field and the induced current at the scatterer [15]. In (14), ${\mathsf{E}}({\bf{r}})$ denotes the total electrical field, which is ${\mathsf{E}}({\bf{r}})={\mathsf{E}}_{\mathrm{i}}({\bf{r}})+{\mathsf{E}}_{\mathrm{s}}({\bf{r}})$, having indicated with ${\mathsf{E}}_{\mathrm{i}}({\bf{r}})$ the electrical field incident to the scatterer generated by the source in the absence of the scatterer and with ${\mathsf{E}}_{\mathrm{s}}({\bf{r}})$ the scattered field caused by the induced current ${\mathsf{J}}_{\mathrm{s}}({\bf{r}})$.
Suppose we aim to solve the EM problem of determining the total electric field ${\mathsf{E}}({\bf{r}})$ and/or its scattered component ${\mathsf{E}}_{\mathrm{s}}({\bf{r}})$ when the scattering device is illuminated by an EM source. Both the electric field and the current must satisfy the free-space propagation equation (1) and the constitutive equation (14), which imposes a boundary condition. Examining the relationship between (1) and (14), it is noteworthy that any scattering device can be viewed as a feedback system, as illustrated in Fig. 4. In this framework, ${\mathsf{E}}_{\mathrm{i}}({\bf{r}})$ represents the input, the propagation operator $\mathbb{G}$ acts as a space-invariant filter, filtering the wavenumber components of the electric field within the support ${\cal{E}}$, as defined in (6), and the scattering device itself functions as a space-variant filter $\mathbb{D}$ (analogous to time-variant filters), with the induced current ${\mathsf{J}}_{\mathrm{s}}({\bf{r}})$ as its output. The filter $\mathbb{D}$ can be designed through the parameters $\bm{\theta}$, i.e., $\mathbb{D}(\bm{\theta})$.
@@
The analysis of the feedback system in Fig. 4 leads to integral equations, the solution of which is generally a complex problem that is typically tackled numerically. A more efficient approach is to use the method of moments, which involves expressing the fields and currents as series expansions using appropriately defined basis sets [37]. To this end, let $\left\{{\mathbf{\Phi}}_{k}({\bf{r}})\right\}_{k=1}^{K}$ denote a complete orthonormal vector basis set for ${\cal{S}}$, which depends solely on the geometry of ${\cal{S}}$. In general, $K$ is infinite for the basis set to be complete, but $K$ can be set to a sufficiently large finite value based on the desired level of accuracy. Consequently, the current density ${\mathsf{J}}_{\mathrm{s}}({\bf{r}})$, the incident and scattered electric fields ${\mathsf{E}}_{\mathrm{i}}({\bf{r}})$, ${\mathsf{E}}_{\mathrm{s}}({\bf{r}})$, and the total electric field ${\mathsf{E}}({\bf{r}})$ within the space ${\cal{S}}$, i.e., for ${\bf{r}}\in{\cal{S}}$, can be expressed as a linear combination of these basis functions, namely
${\mathsf{J}}_{\mathrm{s}}({\bf{r}})\!=\!\sum_{k=1}^{K}[{\bf{j}}_{\mathrm{s}}]_{k}\,{\mathbf{\Phi}}_{k}({\bf{r}})\,,\quad{\mathsf{E}}_{\mathrm{i}}({\bf{r}})\!=\!\sum_{k=1}^{K}[{\bf{e}}_{\mathrm{i}}]_{k}\,{\mathbf{\Phi}}_{k}({\bf{r}})\,,\quad{\mathsf{E}}_{\mathrm{s}}({\bf{r}})\!=\!\sum_{k=1}^{K}[{\bf{e}}_{\mathrm{s}}]_{k}\,{\mathbf{\Phi}}_{k}({\bf{r}})\,,\quad{\mathsf{E}}({\bf{r}})\!=\!\sum_{k=1}^{K}[{\bf{e}}]_{k}\,{\mathbf{\Phi}}_{k}({\bf{r}})$ (15)
@@
where the complex coefficients $\{[{\bf{j}}_{\mathrm{s}}]_{k}\}$ composing the vector ${\bf{j}}_{\mathrm{s}}$ are given by $[{\bf{j}}_{\mathrm{s}}]_{k}=\left<{\mathsf{J}}_{\mathrm{s}}({\bf{r}}),{\mathbf{\Phi}}_{k}({\bf{r}})\right>$, for $k=1,2,\ldots,K$. The vectors ${\bf{e}}_{\mathrm{i}}$, ${\bf{e}}_{\mathrm{s}}$, and ${\bf{e}}$ are defined analogously, containing the coefficients of the series expansions of, respectively, ${\mathsf{E}}_{\mathrm{i}}({\bf{r}})$, ${\mathsf{E}}_{\mathrm{s}}({\bf{r}})$, and ${\mathsf{E}}({\bf{r}})$, according to the same basis set $\{{\mathbf{\Phi}}_{k}({\bf{r}})\}$. The vectors ${\bf{j}}_{\mathrm{s}}$, ${\bf{e}}_{\mathrm{i}}$, ${\bf{e}}_{\mathrm{s}}$, and ${\bf{e}}$ represent generalized currents and voltages. When the scattering device consists of a set of radiating elements (e.g., dipoles), each with a programmable loaded port, a suitable choice of the functions $\left\{{\mathbf{\Phi}}_{k}({\bf{r}})\right\}$ allows ${\bf{j}}_{\mathrm{s}}$ and ${\bf{e}}_{\mathrm{s}}$ to be directly associated with the actual currents and voltages at the ports. This effectively maps the original problem into an electrical circuit model (see Sec. V-A).
@@
From Fig. 4, it straightforward to write ${\mathsf{E}}({\bf{r}})={\mathsf{E}}_{\mathrm{i}}({\bf{r}})+(\mathbb{G}\,{\mathsf{J}}_{\mathrm{s}})({\bf{r}})={\mathsf{E}}_{\mathrm{i}}({\bf{r}})+(\mathbb{G}\,\mathbb{D}\,{\mathsf{E}})({\bf{r}})$. As a consequence, ${\bf{e}}={\bf{e}}_{\mathrm{i}}+{\bf{e}}_{\mathrm{s}}={\bf{e}}_{\mathrm{i}}+{\bf{G}}\,{\bf{j}}_{\mathrm{s}}={\bf{e}}_{\mathrm{i}}+{\bf{G}}\,{\bf{D}}\,{\bf{e}}$, where ${\bf{G}}\in\mathbb{C}^{K\times K}$ is a matrix accounting for the self-coupling of the scattering device having elements $[{\bf{G}}]_{i,k}=\left<{{\mathbf{\Phi}}_{i}({\bf{r}})}\,,{(\mathbb{G}\,{\mathbf{\Phi}}_{k})({\bf{r}})}\right>$ for $i,k=1,2,\ldots,K$, and ${\bf{D}}\in\mathbb{C}^{K\times K}$ is a matrix with elements $[{\bf{D}}]_{i,k}=\left<{{\mathbf{\Phi}}_{i}({\bf{r}})}\,,{(\mathbb{D}\,{\mathbf{\Phi}}_{k})({\bf{r}})}\right>$, for $i,k=1,2,\ldots,K$. It follows that
${\bf{e}}_{\mathrm{s}}={\bf{G}}\,{\bf{j}}_{\mathrm{s}}={\bf{G}}\,{\bf{D}}({\bf{I}}_{K}-{\bf{G}}\,{\bf{D}})^{-1}\,{\bf{e}}_{\mathrm{i}}={\bf{R}}\,{\bf{e}}_{\mathrm{i}}$ (16)

### v3

@@ III-B Communication Modes
For convenience, we introduce the inner product between vector functions $\mathsf{A}({\mathbf{r}})$ and $\mathsf{B}({\mathbf{r}})$, defined on the generic space $\cal{S}$}, as
$\left<{\mathsf{A}({\mathbf{r}})}\,,{\mathsf{B}({\mathbf{r}})}\right>=\int_{{\cal{S}}}\mathsf{A}({\mathbf{r}})\bm{\cdot}\mathsf{B}^{*}({\mathbf{r}})\,\text{d}{\mathbf{r}}\,.$ (11)
Two vector functions are said to be orthogonal if $\left<{\mathsf{A}({\mathbf{r}})}\,,{\mathsf{B}({\mathbf{r}})}\right>=0$.
@@
A fundamental question is determining how many orthogonal communication channels—referred to as communication modes—can be exploited between a source space $\mathcal{S}_{\text{T}}$ and a receiver space $\mathcal{S}_{\text{R}}$ while maximizing the coupling intensity (see Fig. 2). This is associated with the optimal approximation of every element in the image space of a Hilbert-Schmidt operator in terms of singular functions. Specifically, define $\mathcal{X}={\mathbb{L}}^{2}(\mathcal{S}_{\text{T}})$ and $\mathcal{Y}={\mathbb{L}}^{2}(\mathcal{S}_{\text{R}})$ the Hilbert spaces corresponding to the square-integrable functions defined in $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$, respectively. The function ${\mathsf{E}}({\mathbf{r}})\in\mathcal{Y}$ can be seen as the image of ${\mathsf{J}}({\mathbf{r}})\in\mathcal{X}$ through the Hilbert-Schmidt kernel $\mathsf{G}\left({\mathbf{r}},{\mathbf{s}}\right)=\mathsf{G}\left({\mathbf{r}}-{\mathbf{s}}\right)$ on $\mathcal{S}_{\text{T}}\times\mathcal{S}_{\text{R}}$, which induces the operator $\mathbb{G}:\mathcal{X}\rightarrow\mathcal{Y}$ such that, for any ${\mathsf{J}}({\mathbf{r}})\in\mathcal{X}$, we can write (2). Define the following self-adjoint Hilbert-Schmidt operators $\mathbb{G}\,\mathbb{G}^{*}$ and $\mathbb{G}^{*}\,\mathbb{G}$ and let $\{{\bm{\varphi}}_{n}({\mathbf{r}})\}$ and $\{{\mathbf{\Psi}}_{n}({\mathbf{r}})\}$ be the eigenfunctions, associated with eigenvalues $\sigma_{n}^{2}$, of the following coupled eigenfunction problems [37]
$\left(\mathbb{G}\,\mathbb{G}^{*}\,{\bm{\varphi}}_{n}\right)({\mathbf{r}})=\sigma_{n}^{2}\,{\bm{\varphi}}_{n}({\mathbf{r}})\quad\quad\quad\quad\quad\quad\left(\mathbb{G}^{*}\,\mathbb{G}\,{\mathbf{\Psi}}_{n}\right)({\mathbf{r}})=\sigma_{n}^{2}\,{\mathbf{\Psi}}_{n}({\mathbf{r}})\,.$ (12)
@@
An important property of Hilbert-Schmidt operators is that they are compact and admit either a finite or countably infinite orthonormal basis. In particular, two sets of orthonormal eigenfunctions $\{{\bm{\varphi}}_{n}({\mathbf{r}})\}$, $\{{\mathbf{\Psi}}_{n}({\mathbf{r}})\}$ exist, which are solutions, respectively, of the coupled eigenfunction problems in (12) with the same real eigenvalues $\sigma_{1}^{2}\geq\sigma_{2}^{2}\geq\sigma_{3}^{2}\ldots$ [37, 38]. Note that $\left\{{\bm{\varphi}}_{n}({\mathbf{r}})\right\}$ and $\left\{{\mathbf{\Psi}}_{n}({\mathbf{r}})\right\}$ are two sets of orthonormal (vector) functions that are complete, respectively, in $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$. The orthonormal condition implies that $\left<{{\bm{\varphi}}_{k}({\mathbf{r}})}\,,{{\bm{\varphi}}_{n}({\mathbf{r}})}\right>=\delta_{k-n}$, and similarly, the same condition holds for the set $\{{\mathbf{\Psi}}_{n}({\mathbf{r}})\}$. It holds that $\left(\mathbb{G}\,{\bm{\varphi}}_{n}\right)({\mathbf{r}})=\sigma_{n}\,{\mathbf{\Psi}}_{n}({\mathbf{r}})$, that is, a current excitation ${\bm{\varphi}}_{n}({\mathbf{r}})$ in $\mathcal{S}_{\text{T}}$ produces an electrical field in $\mathcal{S}_{\text{R}}$ equal to $\sigma_{n}\,{\mathbf{\Psi}}_{n}({\mathbf{r}})$. In general, the eigenvalues have significant magnitudes up to a certain value $\sigma_{N}^{2}$, i.e., the sum of the $\sigma_{n}^{2}$’s for $n>N$ is negligible with respect to the sum of the first $N$ eigenvalues, so that we can consider up to $N$ well-coupled communication modes between the two spaces. The number $N$ is often called the number of DoF available in the link. Detailed discussion on optimal approximation and the definition of DoF can be found in [37].
@@
In summary, the finite coupling operator $\mathbb{G}$ in (2) between finite spaces $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$ is a Hilbert-Schmidt operator, and up to $N$ communication modes (i.e., orthogonal channels) can be identified for a given level of approximation accuracy, each of them represented by the triad $\left({\bm{\varphi}}_{n}({\mathbf{r}}),{\mathbf{\Psi}}_{n}({\mathbf{r}}),\sigma_{n}\right)$. It follows that the current density ${\mathsf{J}}({\mathbf{r}})$ on the space $\mathcal{S}_{\text{T}}$, i.e., for ${\mathbf{r}}\in\mathcal{S}_{\text{T}}$, can be well approximated as a linear combination of the basis functions
${\mathsf{J}}({\mathbf{r}})$ $\simeq\sum_{n=1}^{N}j_{n}\,{\bm{\varphi}}_{n}({\mathbf{r}})\,$, (13)
@@
where the complex coefficients $j_{n}$ are given by $j_{n}=\left<{\mathsf{J}}({\mathbf{r}}),{\bm{\varphi}}_{n}({\mathbf{r}})\right$>, for $n=1,2,\ldots,N$. Similarly, the electric field ${\mathsf{E}}({\mathbf{r}})$ in $\mathcal{S}_{\text{R}}$ can be approximated through the basis set $\left\{{\mathbf{\Psi}}_{n}({\mathbf{r}})\right\}$ as [38]
${\mathsf{E}}({\mathbf{r}})$ $\simeq\sum_{n=1}^{N}e_{n}\,{\mathbf{\Psi}}_{n}({\mathbf{r}})=\sum_{n=1}^{N}\sigma_{n}\,j_{n}\,{\mathbf{\Psi}}_{n}({\mathbf{r}})\,.$ (14)
It is worth noting that, while in general a basis set for all functions defined within a bounded domain, depending solely on its geometry, can always be identified (for instance, the harmonic functions of the Fourier series [15]), the optimal approximating basis sets $\left({\bm{\varphi}}_{n}({\mathbf{r}}),{\mathbf{\Psi}}_{n}({\mathbf{r}})\right)$ obtained as the solution of (12) also depend on the surrounding environment through the Green’s coupling operator $\mathbb{G}$, in addition to $\mathcal{S}_{\text{T}}$ and $\mathcal{S}_{\text{R}}$.
@@
Figure 3: A source and receiver space interacting through a reconfigurable scattering device acting as an EM signal processor.
@@ III-C Reconfigurable Scattering Device: Mode Transfer Matrix
Consider now the scenario illustrated in Fig. 3, where a source space $\mathcal{S}_{\text{T}}$ (e.g., a transmitter) and a receiver space $\mathcal{S}_{\text{R}}$ (the receiver) interact through a reconfigurable scattering device. We assume the scattering device to be enclosed within the finite space $\cal{S}$}. The impinging EM field generated by currents located in the source space $\mathcal{S}_{\text{T}}$ induces a current distribution ${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})$ on the scattering device, with ${\mathbf{r}}\in{\cal{S}}$, that radiates the scattered field ${\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})$ [35]. [^4] A more general treatment is based on the equivalent principle where fictitious electric and magnetic currents are introduced as a function of the tangent electric and magnetic fields [15]. For clarity, hereby we focus on a more limited case valid for dielectric and non-perfect conductors.
For any linear time-invariant scattering object, the induced current density ${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})$, ${\mathbf{r}}\in{\cal{S}}$, is a linear functional of the EM field ${\mathsf{E}}({\mathbf{r}})$ that can be described as
${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})=\int_{{\cal{S}}}\mathsf{D}({\mathbf{r}},{\mathbf{s}})\,{\mathsf{E}}({\mathbf{s}})\,d{\mathbf{s}}=\left(\mathbb{D}\,{\mathsf{E}}\right)({\mathbf{r}})\,$, (15)
where $\mathsf{D}({\mathbf{r}},{\mathbf{s}})$ is the dyadic impulse response that fully describes the relationship between the electric field and the induced current at the scatterer [15]. In (15), ${\mathsf{E}}({\mathbf{r}})$ denotes the total electrical field, which is ${\mathsf{E}}({\mathbf{r}})={\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})+{\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})$, having indicated with ${\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})$ the electrical field incident to the scatterer generated by the source in the absence of the scatterer and with ${\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})$ the scattered field caused by the induced current ${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})$.
@@
The analysis of the feedback system in Fig. 4 leads to integral equations, the solution of which is generally a complex problem that is typically tackled numerically. A more efficient approach is to use the method of moments, which involves expressing the fields and currents as series expansions using appropriately defined basis sets [35]. To this end, let $\left\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\right\}_{k=1}^{K}$ denote a complete orthonormal vector basis set for $\cal{S}$}, which depends solely on the geometry of $\cal{S}$}. In general, $K$ is infinite for the basis set to be complete, but $K$ can be set to a sufficiently large finite value based on the desired level of accuracy. Consequently, the current density ${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})$, the incident and scattered electric fields ${\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})$, ${\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})$, and the total electric field ${\mathsf{E}}({\mathbf{r}})$ within the space $\cal{S}$}, i.e., for ${\mathbf{r}}\in{\cal{S}}$, can be expressed as a linear combination of these basis functions, namely
${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})\!=\!\sum_{k=1}^{K}[{\mathbf{j}}_{\mathrm{s}}]_{k}\,{\mathbf{\Phi}}_{k}({\mathbf{r}})\,,\quad{\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})\!=\!\sum_{k=1}^{K}[{\mathbf{e}}_{\mathrm{i}}]_{k}\,{\mathbf{\Phi}}_{k}({\mathbf{r}})\,,\quad{\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})\!=\!\sum_{k=1}^{K}[{\mathbf{e}}_{\mathrm{s}}]_{k}\,{\mathbf{\Phi}}_{k}({\mathbf{r}})\,,\quad{\mathsf{E}}({\mathbf{r}})\!=\!\sum_{k=1}^{K}[{\mathbf{e}}]_{k}\,{\mathbf{\Phi}}_{k}({\mathbf{r}})\,$, (16)
@@
where the vectors ${\mathbf{j}}_{\mathrm{s}}$, ${\mathbf{e}}_{\mathrm{i}}$, ${\mathbf{e}}_{\mathrm{s}}$, and $\mathbf{e}$} represent generalized currents and voltages. The complex coefficients $\{[{\mathbf{j}}_{\mathrm{s}}]_{k}\}$ composing the vector ${\mathbf{j}}_{\mathrm{s}}$ are given by ${\mathbf{j}}_{\mathrm{s}}]_{k}=\left<{\mathsf{J}}_{\mathrm{s}}({\mathbf{r}}),{\mathbf{\Phi}}_{k}({\mathbf{r}})\right$>, for $k=1,2,\ldots,K$. The vectors ${\mathbf{e}}_{\mathrm{i}}$, ${\mathbf{e}}_{\mathrm{s}}$, and $\mathbf{e}$} are defined analogously, containing the coefficients of the series expansions of, respectively, ${\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})$, ${\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})$, and ${\mathsf{E}}({\mathbf{r}})$, according to the same basis set $\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\}$. When the scattering device consists of a set of radiating elements (e.g., dipoles), each with a programmable loaded port, a suitable choice of the functions $\left\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\right\}$ allows ${\mathbf{j}}_{\mathrm{s}}$ and ${\mathbf{e}}_{\mathrm{s}}$ to be directly associated with the actual currents and voltages at the ports. This effectively maps the original problem into an electrical circuit model (see Section V-A).
@@
From Fig. 4, it straightforward to write ${\mathsf{E}}({\mathbf{r}})={\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})+(\mathbb{G}\,{\mathsf{J}}_{\mathrm{s}})({\mathbf{r}})={\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})+(\mathbb{G}\,\mathbb{D}\,{\mathsf{E}})({\mathbf{r}})$. As a consequence, ${\mathbf{e}}={\mathbf{e}}_{\mathrm{i}}+{\mathbf{e}}_{\mathrm{s}}={\mathbf{e}}_{\mathrm{i}}+{\mathbf{G}}\,{\mathbf{j}}_{\mathrm{s}}={\mathbf{e}}_{\mathrm{i}}+{\mathbf{G}}\,{\mathbf{D}}(\bm{\theta})\,{\mathbf{e}}$, where ${\mathbf{G}}\in\mathbb{C}^{K\times K}$ is a matrix accounting for the self-coupling of the scattering device having elements ${\mathbf{G}}]_{i,k}=\left<{{\mathbf{\Phi}}_{i}({\mathbf{r}})}\,,{(\mathbb{G}\,{\mathbf{\Phi}}_{k})({\mathbf{r}})}\right$> for $i,k=1,2,\ldots,K$, and ${\mathbf{D}}(\bm{\theta})\in\mathbb{C}^{K\times K}$ is a matrix with elements ${\mathbf{D}}(\bm{\theta})]_{i,k}=\left<{{\mathbf{\Phi}}_{i}({\mathbf{r}})}\,,{(\mathbb{D}\,{\mathbf{\Phi}}_{k})({\mathbf{r}})}\right$>, for $i,k=1,2,\ldots,K$. It follows that
${\mathbf{e}}_{\mathrm{s}}={\mathbf{G}}\,{\mathbf{j}}_{\mathrm{s}}={\mathbf{G}}\,{\mathbf{D}}(\bm{\theta})({\mathbf{I}}_{K}-{\mathbf{G}}\,{\mathbf{D}}(\bm{\theta}))^{-1}\,{\mathbf{e}}_{\mathrm{i}}={\mathbf{R}}(\bm{\theta})\,{\mathbf{e}}_{\mathrm{i}}\,$, (17)